Tina drove from Calgary to Vancouver, a distance of 1000km in exactly 12 hours 18 minutes. She drove at a speed of 100km/h on the highway and 50km/h in the towns. What distance did she drive at each speed?
Answer by mananth(16946)

Let x be the speed at 100 km/h
y be speed at 50 km/h
total distance = 1000 km
d=rt
100x+50y = 1000
time
x+y = 12.3
multiply bu 100
100x + 100y = 1230
subtract the equations0y = 230
y= 4.6
distance at 50 km/h = 4.6 * 50 = 230km
7.7 hours at 100 km/h = 770 km